# Figure 1. Confirmation–Diagnosis–Correction–Redo (CDCR) Pattern. Based on the expected time of user interaction under the CDCR pattern, our model decides whether the agent should proceed autonomously or prompt the user to step in and supervise.:

### v1

@@ Abstract.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
With the advancement of Artificial Intelligence (AI), Agentic AI has emerged as a concept describing agents that assist humans in carrying out complex tasks (Durante et al., 2024; Pallagani et al., 2024; Huang et al., 2024; Wang et al., 2024a; Yao et al., 2023; Sumers et al., 2023; Schick et al., 2023; Wei et al., 2022). These AI agents have been deployed across diverse domains, including multi-modal devices (rabbit research team, 2023; Chang et al., 2024), web-based systems (Huq et al., 2023; Li et al., 2023, 2017), and game controls (Wang et al., 2023b, a; Lawley and MacLellan, 2024). Their rapid proliferation has reignited a long-standing HCI debate (Horvitz, 1999; Shneiderman and Maes, 1997; Lieberman, 1997; Maes, 1995): should agents operate under direct human supervision and control or should they operate in a mostly autonomous fashion?
Most commercial AI agents today prioritize autonomy, executing tasks end-to-end with little to no user oversight (OpenAI, 2025; Anthropic, 2024; Manus, 2024). Yet these agents remain highly error-prone, achieving only  30% accuracy on multi-step benchmarks (Ma et al., 2024; Xie et al., 2024a; Xu et al., 2024), largely due to their limited capacity for generalizable reasoning and planning (Stechly et al., 2024; Valmeekam et al., 2024). Errors are especially costly in long-horizon agentic tasks, which often span 30 steps and 10–30 minutes of execution (Hu et al., 2024; Bonatti et al., 2024). A single mistake can cascade into complete task failure, forcing users to redo the entire process—doubling execution time and, through repeated API calls, adding both monetary costs (Kapoor et al., 2024; Feng et al., 2024; Zhang et al., 2025a) and increasing the agent’s carbon footprint (Patterson et al., 2022; Luccioni et al., 2023). This situation highlights the need to engage user inspections over agent execution.
@@
Prior HCI research has largely focused on how to provide user control, through visualizing the reasoning process, editing execution results, and providing error feedback (Arawjo et al., 2024; Epperson et al., 2025; Kazemitabaar et al., 2024; Feng et al., 2024; Lam et al., 2022; Ashktorab et al., 2019; Li et al., 2020; Levy et al., 2021). Few studies examine when such control should occur. A notable exception is Horvitz’s seminal work on expected utility models for single-event decisions (Horvitz, 1999), but these approaches do not extend to long-horizon tasks, where sequential dependencies and cumulative costs shape overall outcomes.
In this paper, we address the open question of when user checks should occur in long-horizon agentic tasks. We first review representative agent benchmarks and deployed systems and present the results of a formative study we conducted with eight participants. The study revealed dissatisfaction with the confirm-at-end strategy used by existing agentic platforms. It also surfaced a recurring Confirmation–Diagnosis–Correction–Redo (CDCR) pattern employed by users to supervise agents. Building on these insights, we developed a decision-theoretic model that identifies periodic confirmation points that minimize total expected task completion time while ensuring task correctness. Finally, we directly validated our approach by comparing it to the confirm-at-end strategy using a within-subjects study with 48 participants. Results show that 81% of participants (39 of 48) preferred our intermediate confirmation strategy over the confirm-at-end baseline. We also found that our strategy reduced the average task completion time by 13.54% compared to the baseline.
@@ 2.1. Human Control in AI Agents
The tension between system autonomy and direct user control has been a recurring theme in HCI for decades (Horvitz, 1999; Shneiderman and Maes, 1997; Lieberman, 1997; Maes, 1995). Current AI agents are highly automated, integrating context, tools, and memory to solve complex real-world tasks (Thrun, 2002). Yet these reasoning and tool use systems are highly error-prone, yielding only  30% accuracy on 10-step tasks (Wei et al., 2022; Dziri et al., 2023). Errors propagate like a snowball—growing exponentially and causing downstream failures (Zhao et al., 2024b; Ji et al., 2023). Agent self-correction is inconsistent and does not scale with task complexity (Kadavath et al., 2022). These challenges highlight the necessity of human control.
@@
To support such control, HCI researchers have proposed a wide range of tools for human involvement in agent planning and execution (Shao et al., 2024; Das Antar et al., 2024; Zhao et al., 2024a; Suh et al., 2024; Feng et al., 2024; Lawley and MacLellan, 2024; Epperson et al., 2025; Kazemitabaar et al., 2024). For instance, Cocoa builds on computational notebook interfaces to support co-planning and co-execution of research tasks (Feng et al., 2024), AGDebugger enables pausing and editing agent behaviors in programming (Epperson et al., 2025), VAL lets users instruct agents to execute tasks in user-specified ways (Lawley and MacLellan, 2024), and Step-Phasewise decomposes complex analysis workflows for verification (Kazemitabaar et al., 2024). These systems focus on how to give users control, but not when that control should be invoked in multi-step tasks.
@@
If we want to study ”when”, we inevitably need mathematical models. Beyond the agentic AI context, prior work has examined the timing of user engagement through mathematical models (Horvitz, 1999; Cockburn et al., 2022; Todi et al., 2021; Zhao et al., 2022). For instance, Horvitz (1999) proposed expected-utility models for deciding whether to invoke services, and Cockburn et al. (2022) showed that users’ probability-weighting biases shape when they engage or ignore assistive features. However, these works focus on single-event decisions and are not applicable to agentic AI, where tasks involve many interdependent steps. To capture the long-horizon nature of agent execution, we need to observe users’ sequential behaviors around errors and build mathematical models that identify the optimal moments for user confirmation from a global, system-level perspective.
@@ 2.2. Error Prevention and Recovery Models
Figure 2. Representative Task Types and AI Agent Interface Example:
@@ 3.1.1. Task Types
To understand the landscape of agentic AI tasks and guide the design of our user study, we reviewed 12 benchmarks that cover a wide spectrum of real-world scenarios (Xie et al., 2024a; Valmeekam et al., 2023; Xie et al., 2024b; Zhou et al., 2023; Xu et al., 2024; Ma et al., 2024; Jing et al., 2024; Yao et al., 2022; Wei et al., 2025; Wang et al., 2024b; Chang et al., 2024). These benchmarks are grounded in user-generated platforms such as StackOverflow, Reddit, and other public-facing systems, where users frequently ask how to complete complex or ambiguous tasks. Each benchmark contains around 300 pieces of task data. From these benchmarks, we identified four broad application domains—Office, Daily Life, Virtual Environment, and Mixed Workflow—each encompassing several representative task types (Figure 2, left). These domains provided the basis for selecting tasks in our formative study. Despite advances in LLMs, current models such as GPT-4 or Claude 3.5 still struggle to reliably complete long-horizon tasks. Reported success rates for benchmarks like OSWorld (Xie et al., 2024a), SpreadsheetBench (Ma et al., 2024), and TheAgentCompany (Xu et al., 2024) often fall between 30% and 60%, with failure modes including incorrect tool use, misinterpreted instructions, or skipped steps. This relatively high error rate underscores the need for human-in-the-loop oversight, where users monitor and intervene during task execution.
@@ 3.1.2. Existing Systems
From the interaction perspective, we observed three common UI patterns across existing agents: (1) in some agents, each step is a clickable node linked to its corresponding screenshot (Manus, 2024; Flowith, 2024; CrewAI, 2024; AI, 2024a); (2) in others, the agent’s activity or reasoning process is visually embedded within the screenshot, with a sliding progress bar to navigate through the log (OpenAI, 2025; Anthropic, 2024; Genspark, 2024); and (3) some agents provide only screenshot playback without an activity log (AI, 2024b; Use, 2024). These UI patterns focus only on passively displaying the agent’s execution, but they do not proactively engage users to verify correctness. Given the low accuracy rates of current agents and the high cost of errors, confirmation mechanisms are exceedingly rare: most systems execute entire task sequences autonomously and pause only when they require missing information (e.g., payment credentials) (OpenAI, 2025). Most errors (e.g., failed login, page not loading) are not detected in real time, and agents simply stop without notifying the user or recovering (Kadavath et al., 2022). These gaps motivated us to explore how to schedule intermediate confirmations in ways that ensure timely user intervention without introducing unnecessary overhead.
@@ 3.2.1. Environment Setup
This diversity of agents and tasks allows us to examine whether consistent patterns of user supervision behavior emerge across different UI designs and task types.
@@ 4. Formative Study Findings
Our thematic analysis of the formative study data identified a dominant Confirmation–Diagnosis–Correction–Redo (CDCR) pattern where participants dealt with AI agent errors by reviewing the execution history from beginning to end. This strategy was evident across diverse task types and UI designs. We also observed some exceptions to this pattern (RQ1). For RQ2, participants expressed a desire for confirmation strategies beyond the prevalent “confirm-at-the-end” norm, suggesting opportunities for intermediate checkpoints. These insights informed the framing of our confirmation frequency model (section 4.3).
@@ 4.1. Confirm-at-end is not enough
These findings suggest the dominant “confirm-at-the-end” approach should be reconsidered. Introducing intermediate confirmations can improve efficiency by allowing users to ensure the agent stays on the right track. Confirming every step is impractical: “If I have to check every step my agent is making, I’d rather do it myself. (P6)” The challenge, therefore, is to strike a balance between the fixed time cost of reviewing every agent action and the expected time cost of needing to confirm task completion at the end and then possibly undertake a long post-hoc diagnostic and redo process if the agent failed. We discuss our model for balancing between these two extremes with checkpoint scheduling in the next section.
@@ 4.2.3. Exceptions to the CDCR pattern
We found that interface differences did not fundamentally change user behavior. However, variations in task type and user familiarity revealed several edge cases where this pattern did not hold. (1) Ambiguous confirmability: In some tasks, there was no clear ground truth for what counted as an error. For example, in hotel booking, the agent might choose a location that was slightly more expensive but offered better transportation access. While arguably suboptimal, this choice was not objectively wrong. (2) Non-linear diagnosis: Some users skipped intermediate steps during diagnosis. For example, P2, who was highly experienced with Overcooked, noticed the player placing soup on the wrong pad and bypassed the cooking phase, jumping directly to the serving phase. (3) Partial correction: When errors were minor and unlikely to affect subsequent planning, users preferred to download the result file and fix the error themselves rather than having the agent re-execute the task.
@@ 4.3. Implications for Model Scope and Applicability
Building on the recurring confirmation–diagnosis–correction–redo pattern, our model is grounded in three key assumptions:
@@ 5.1.1. State space $\mathcal{S}$
where $1<i\leq N$ and $p_{a_{i}}$ corresponds to the probability that the agent will execute action $i$ from its plan correctly. These transitions also capture the cumulative probability of errors; i.e., once an agent makes a mistake, all subsequent states will be incorrect. Note that there is no execute transition out of the $s_{N}$ state. An important point here is that the correctness of all states after the last user verified state is not directly observable to the system; it becomes known only when the user inspects them. To compute the probability that any given state $s_{i}$ is correct, the system must weigh the transition from the previous state by its beliefs about the correctness of the prior state: $p(s_{i}=\texttt{correct})=p(s_{i-1}=\texttt{correct})p(s_{i}=\texttt{correct}|s_{i-1}=\texttt{correct},\texttt{execute})$. By chaining over transitions, we can compute the probability that any state $j$ is correct given the last user verified state $i$, where $i<j$: $p(s_{j}=\texttt{correct}|s_{i}=\texttt{correct})=\prod_{m=i+1}^{j}p_{a_{m}}$. Note, we do not need to consider actions before the last user verified step because we already know they are all correct.
@@
Figure 4. Dynamic programming algorithm and results for an agent action sequence of length $N=5$ with action success probabilities $p=[0.7,0.7,0.9,0.85,0.85]$. Costs are parameterized as $t_{\text{confirm}}[k]=1$, $t_{\text{diagnose}}[k]=1$, and $t_{\text{redo}}[k]=1$, for all $k\in\{0,\dots,N\}$. Each cell $\textit{start},\textit{end}]$ specifies the expected user time to correctly execute the rest of the action sequence from startstart to $N$, if the system has a verification checkpoint at state end. White checkmarks indicate the next_ckptnext\_ckpt locations for each startstart. The right-hand annotations summarize the resulting checkpoint policy;

### v2

# Confirmation–Diagnosis–Correction–Redo (CDCR)

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790655isbn: 979-8-4007-2278-3/2026/04
@@ 1. Introduction
With the advancement of Artificial Intelligence (AI), Agentic AI has emerged as a concept describing agents that assist humans in carrying out complex tasks (Durante et al., 2024; Pallagani et al., 2024; Huang et al., 2024; Wang et al., 2024a; Yao et al., 2023; Sumers et al., 2023; Schick et al., 2023; Wei et al., 2022). These AI agents have been deployed across diverse domains, including multi-modal devices (rabbit research team, 2023; Chang et al., 2024), web-based systems (Huq et al., 2023; Li et al., 2023, 2017), and game controls (Wang et al., 2023a, b; Lawley and MacLellan, 2024). Their rapid proliferation has reignited a long-standing HCI debate (Horvitz, 1999; Shneiderman and Maes, 1997; Lieberman, 1997; Maes, 1995): should agents operate under direct human supervision and control or should they operate in a mostly autonomous fashion?
Most commercial AI agents today prioritize end-to-end autonomy, and rarely proactively ask users to confirm intermediate steps (OpenAI, 2025; Anthropic, 2024; Manus, 2024). While some agents allow users to take control or interrupt execution (OpenAI, 2025; Qin et al., 2025), these mechanisms are reactive: they are only initiated when users halt execution upon noticing something is wrong, or when the agent halts to request missing information. Without proactive reminders, users often miss the window for timely intervention. As a result, errors frequently go unnoticed until the end of the trajectory. This is problematic because current agents remain highly error-prone, achieving only  30% accuracy on multi-step benchmarks (Ma et al., 2024; Xie et al., 2024b; Xu et al., 2024), largely due to their limited capacity for generalizable reasoning and planning (Stechly et al., 2024; Valmeekam et al., 2024). Errors are especially costly in long-horizon agentic tasks, which often span 30 steps and 10–30 minutes of execution (Hu et al., 2024; Bonatti et al., 2024). A single mistake can cascade into complete task failure, forcing users to redo the entire process—doubling execution time and, through repeated API calls, adding both monetary costs (Kapoor et al., 2024; Feng et al., 2024; Zhang et al., 2025b) and increasing the agent’s carbon footprint (Patterson et al., 2022; Luccioni et al., 2023). This motivates the need to determine when agents should proactively elicit user confirmation to prevent cascading failures.
@@
Prior HCI research has largely focused on how to provide user control, through visualizing the reasoning process, editing execution results, and providing error feedback (Arawjo et al., 2024; Epperson et al., 2025; Kazemitabaar et al., 2024; Feng et al., 2024; Lam et al., 2022; Ashktorab et al., 2019; Li et al., 2020; Levy et al., 2021). Some recent systems also consider when assistance should occur in user-led tasks, where the human is actively performing the activity and the agent decides when to offer help based on heuristic triggers or models of the user’s internal state (Pu et al., 2025b; Li et al., 2025a; Arakawa et al., 2024; Pu et al., 2025a). In contrast, we study the mirror case of agent-led workflows: an automated agent executes a multi-step task on the user’s behalf. The key question in this case is when the user should be brought back in to confirm. A notable exception is Horvitz’s seminal work on expected utility models for single-event decisions (Horvitz, 1999), but these approaches do not extend to long-horizon tasks, where sequential dependencies and cumulative costs shape overall outcomes.
In this paper, we address the open question of when user checks should occur in long-horizon agentic tasks. We first review representative agent benchmarks and deployed systems and present the results of a formative study we conducted with eight participants. The study revealed dissatisfaction with the confirm-at-end strategy used by existing agentic platforms. It also surfaced a recurring Confirmation–Diagnosis–Correction–Redo (CDCR) pattern employed by users to supervise agents. Building on these insights, we developed a decision-theoretic model that identifies periodic confirmation points that minimize total expected task completion time while ensuring task correctness. We directly validated our approach by comparing it to the confirm-at-end strategy using a within-subjects study with 48 participants under three domains. Results show that 81% of participants (39 of 48) preferred our intermediate confirmation strategy over the confirm-at-end baseline. We also found that our strategy reduced the average task completion time by 13.54% compared to the baseline. We further discuss parameter realism, present simulations that demonstrate the model’s behavior under step-wise heterogeneity, and outline how our algorithm can be integrated to existing agent systems as a lightweight timing layer that triggers user-control hooks.
@@ 2.1. Human Control in AI Agents
The tension between system autonomy and direct user control has been a recurring theme in HCI for decades (Horvitz, 1999; Shneiderman and Maes, 1997; Lieberman, 1997; Maes, 1995). Current AI agents are highly automated, integrating context, tools, and memory to solve complex real-world tasks (Thrun, 2002). Yet these reasoning and tool-use systems are highly error-prone, yielding only  30% accuracy on 10-step tasks (Wei et al., 2022; Dziri et al., 2023). Errors propagate like a snowball—growing exponentially and causing downstream failures (Zhao et al., 2024a; Ji et al., 2023). Agent self-correction is inconsistent and does not scale with task complexity (Kadavath et al., 2022). These challenges highlight the necessity of human control.
Along this mixed-initiative spectrum, recent work has examined user-led tasks in which the human is the primary actor and the system decides when to offer assistance. Systems such as Satori, PrISM-Observer, ProMemAssist, and proactive programming assistants model users’ activity or cognitive state and trigger help at predicted good moments (Li et al., 2025a; Arakawa et al., 2024; Pu et al., 2025b, a). For example, agents intervene when a user is likely to reach a particular step, forget an action, or benefit from a reminder or suggestion. In these systems, the core timing question is when an assistant should interrupt or scaffold an ongoing user task.
@@
On the agent-led side, HCI researchers have proposed a wide range of tools for human involvement in domain-specific workflows, particularly in programming and data analysis (Shao et al., 2024; Das Antar et al., 2024; Zhao et al., 2024b; Suh et al., 2024; Feng et al., 2024; Lawley and MacLellan, 2024; Epperson et al., 2025; Kazemitabaar et al., 2024; Chen et al., 2023; Zhou, 2025). For instance, Cocoa builds on computational notebook interfaces to support co-planning and co-execution of research tasks (Feng et al., 2024), AGDebugger enables pausing and editing agent behaviors in programming (Epperson et al., 2025), and Step-Phasewise decomposes complex analysis workflows for verification (Kazemitabaar et al., 2024). In addition, research has expanded to general-purpose agents operating in open-ended environments. Systems like CowPilot (Huq et al., 2025) and Magentic-UI (Mozannar et al., 2025) have introduced mixed-initiative controls, enabling users to pause execution, reject plans, or approve specific steps. Similarly, Morae (Peng et al., 2025) and recent image creation agents (Hahn et al., 2025) focus on detecting ambiguous decision points or maintaining explicit intent states to solicit user clarifications dynamically. These systems focus on how to give users control, but not when that control should be invoked in multi-step tasks. Our motivation is to provide a scheduling layer that triggers those existing hooks.
@@
One useful way to study when interventions occur in agent-led workflows is through computational models of timing (Horvitz, 1999; Cockburn et al., 2022; Todi et al., 2021; Zhao et al., 2022). For instance, Horvitz (1999) proposed expected-utility models for deciding whether to invoke services, and Cockburn et al. (2022) showed that users’ probability-weighting biases shape when they engage or ignore assistive features. However, these works focus on single-event decisions and are not applicable to agentic AI, where tasks involve many interdependent steps. To capture the long-horizon nature of agent execution, we need to observe users’ sequential behaviors around errors and build mathematical models that identify the optimal moments for user confirmation from a global, system-level perspective.
@@ 2.2. Error Prevention and Recovery Models
Figure 2. Representative Task Types and AI Agent Interface Example
Left: Task taxonomy generated from 12 representative benchmarks. Right: Example of an AI agent interface.
@@ 3.1.1. Task Types
To understand the landscape of agentic AI tasks and guide the design of our user study, we reviewed 12 benchmarks that cover a wide spectrum of real-world scenarios (Xie et al., 2024b; Valmeekam et al., 2023; Xie et al., 2024a; Zhou et al., 2023; Xu et al., 2024; Ma et al., 2024; Jing et al., 2024; Yao et al., 2022; Wei et al., 2025; Wang et al., 2024b; Chang et al., 2024). These benchmarks are grounded in user-generated platforms such as StackOverflow, Reddit, and other public-facing systems, where users frequently ask how to complete complex or ambiguous tasks. Each benchmark contains around 300 pieces of task data. From these benchmarks, we identified four broad application domains—Office, Daily Life, Virtual Environment, and Mixed Workflow—each encompassing several representative task types (Figure 2, left). These domains provided the basis for selecting tasks in our formative study. Despite advances in LLMs, current models such as GPT-4 or Claude 3.5 still struggle to reliably complete long-horizon tasks. Reported success rates for benchmarks like OSWorld (Xie et al., 2024b), SpreadsheetBench (Ma et al., 2024), and TheAgentCompany (Xu et al., 2024) often fall between 30% and 60%, with failure modes including incorrect tool use, misinterpreted instructions, or skipped steps. This relatively high error rate underscores the need for human-in-the-loop oversight, where users monitor and intervene during task execution.
@@ 3.1.2. Existing Systems
From the interaction perspective, we observed three common UI patterns across existing agents: (1) in some agents, each step is a clickable node linked to its corresponding screenshot (Manus, 2024; Flowith, 2024; CrewAI, 2024; AI, 2024a); (2) in others, the agent’s activity or reasoning process is visually embedded within the screenshot, with a sliding progress bar to navigate through the log (OpenAI, 2025; Anthropic, 2024; Genspark, 2024); and (3) some agents provide only screenshot playback without an activity log (AI, 2024b; Use, 2024). While these UI patterns allow users to inspect what the agent has done, confirmation mechanisms are exceedingly rare, given the low accuracy rates of current agents and the high cost of errors. Most systems execute entire task sequences autonomously, while some can pause to let users take control (OpenAI, 2025; Qin et al., 2025). However, this pause mechanism is passive and reactive: the pause occurs only because the agent cannot proceed (e.g., missing payment credentials or login details), not because it seeks confirmation from users whether its execution so far is correct. As a result, reactive control provides no guidance on when users should check the agent’s progress, because it is entirely user-initiated. Consequently, most errors (e.g., clicking the wrong UI element or failing to load a page) are not detected in real time, and the agent either continues executing incorrectly or stops without recovery (Kadavath et al., 2022). These gaps motivated us to explore how to schedule intermediate confirmations in ways that ensure timely user intervention without introducing unnecessary overhead.
@@ 3.2.1. Environment Setup
The Overcooked task runs in a custom game environment that requires direct programmatic control and access to internal state. In our setup, only Browser Use could reliably operate in this environment, so the Overcooked task was always paired with Agent 3. The remaining four tasks all use conventional web-based interfaces and are similar in complexity, so we randomly assigned them to Agent 1 and Agent 2 while ensuring that each handled one everyday task and one office workflow. This design gives us coverage of three distinct UI patterns while keeping task–agent pairing a nuisance factor rather than a primary variable of interest.
@@ 4. Formative Study Findings
Our thematic analysis of the formative study data identified a dominant Confirmation–Diagnosis–Correction–Redo (CDCR) pattern where participants dealt with AI agent errors by reviewing the execution history from beginning to end. This strategy was evident across diverse task types and UI designs. We also observed some exceptions to this pattern (RQ1). For RQ2, participants expressed a desire for confirmation strategies beyond the prevalent “confirm-at-end” norm, suggesting opportunities for intermediate checkpoints. These insights informed the framing of our confirmation frequency model.
@@ 4.1. Confirm-at-end is not enough
These findings suggest the dominant “confirm-at-the-end” approach should be reconsidered. Although Agent 2 (OpenAI, 2025) offers a reactive control mode (allowing users to seize control of the agent’s screen), six of our eight participants reported that reactive control still made them uncertain about when to check, failing to help them “spend less time monitoring” (P1). Current agent UIs left users in a passive “wait state” during execution, with no meaningful hooks to monitor or steer the run. Consequently, to catch errors in real time, users are forced to maintain constant vigilance; if they look away, errors are easily missed. In that case, users cannot disengage while the agent executes, which undermines the value of automation. As P6 noted, “If I have to check every step my agent is making, I’d rather do it myself.” These results highlight the need for a proactive, agent-initiated confirmation mechanism that reduces users’ monitoring burden. The central challenge is to strike a balance between the fixed time cost of reviewing every agent action and the expected time cost of needing to confirm task completion at the end and then possibly undertake a long post-hoc diagnostic and redo process if the agent failed. We discuss our model for balancing between these two extremes with checkpoint scheduling in the next section.
@@ 4.2.3. Exceptions to the CDCR pattern
Across all 40 trials (5 tasks $\times$ 8 participants), 100% (40/40) of trials entered the Confirmation stage, 87.5% (35/40) proceeded to Diagnosis, and 82.5% (33/40) executed a Correction. Agent differences did not fundamentally change this behavior. However, variations in task type and user familiarity revealed several edge cases in how the later CDCR stages were instantiated. (1) Ambiguous confirmability: In 5 trials (2 hotel-booking and 3 Overcooked), participants failed to identify an existing error. In the hotel-booking domain, ambiguity in what counted as “wrong” made it difficult to diagnose issues: for example, the agent sometimes chose a hotel that was slightly more expensive but had better transportation access. While arguably suboptimal, this choice was not objectively wrong. In the three Overcooked trials, participants noticed the unexpected result and entered the diagnosis phase, but failed to identify the first error. Because the agent continued to act after the initial mistake, participants reported that they expected the agent to correct itself later, and therefore overlooked the original error location. (2) Non-linear diagnosis: Two users skipped intermediate steps during diagnosis in Overcooked. Because they are highly experienced with Overcooked, they jumped directly to key states: after noticing the player placing soup on the wrong pad. (3) Partial correction: In 2 document-processing trials, participants entered the confirmation and diagnosis stages but chose an alternative correction path. They preferred to download the result file and fix the error themselves rather than having the agent re-execute the task, because the errors were minor and unlikely to affect subsequent planning.
@@ 4.3. Implications for Model Scope and Applicability
Edge cases in the previous subsection provide concrete examples of how users actually confirm, diagnose, and correct agent errors in practice. Building on these observations, we abstract a set of modeling assumptions that capture a common regime while making clear where the model does not apply.
@@
  4. (4)
    Once the current step is verified, the model treats all earlier steps as correct and only searches for errors in later steps.
@@ 5.1.1. State space $\mathcal{S}$
where $1<i\leq N$ and $p_{a_{i}}$ corresponds to the probability that the agent will execute action $i$ from its plan correctly. These transitions also capture the cumulative probability of errors; i.e., once an agent makes a mistake, all subsequent states will be incorrect. Note that there is no execute transition out of the $s_{N}$ state. An important point here is that the correctness of all states after the last user-verified state is not directly observable to the system; it becomes known only when the user inspects them. To compute the probability that any given state $s_{i}$ is correct, the system must weigh the transition from the previous state by its beliefs about the correctness of the prior state: $p(s_{i}=\texttt{correct})=p(s_{i-1}=\texttt{correct})p(s_{i}=\texttt{correct}|s_{i-1}=\texttt{correct},\texttt{execute})$. By chaining over transitions, we can compute the probability that any state $j$ is correct given the last user-verified state $i$, where $i<j$: $p(s_{j}=\texttt{correct}|s_{i}=\texttt{correct})=\prod_{m=i+1}^{j}p_{a_{m}}$. Note, we do not need to consider actions before the last user-verified step because we already know they are all correct.